OpenStudy (anonymous):

1 MEDAL AND FAN QUESTION!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

DFG + JKL = 90 x+5 + (x-9) = 90 2x -4 = 90 2x = 94 x = ? With x = ? find (x+5) = DFG and (x-9) = JKL